The cost of debt :  Gronseth Drywall​ Systems, Inc., is in discussions with its investment bankers regarding the issuance of new bonds. The investment banker has informed the firm that different maturities will carry different coupon rates and sell at different prices. The firm must choose among several alternatives. In each​ case, the bonds will have a ​$1000 par value and flotation costs will be ​$30 per bond. The company is taxed at 23​%. Use the approximation formula to calculate the ​after-tax cost of financing with the following alternative.  ​

Coupon rate Time to maturity Premium or discount

9​% 15 years $220

The​ after-tax cost of financing using the approximation formula is ?  ​(Round to two decimal​ places.)
Answer #1

Premium

B12 - X Fx =B9*(1-B11) B 1000 0.09 =B2*B1 220 =B1+B4 30 1 Par Value 2 Coupon rate 3 Coupon payment 4. Discount 5 Market Value

B12 : x for =B9*(1-B11) B 1000 9% 90 220 1220 30 1 Par Value 2 Coupon rate 3 Coupon payment 4. Discount 5 Market Value 6 Flot

Discount

B12 - X = fx =B9*(1-B11) B 1000 9% 90 220 1 Par Value 2 Coupon rate 3 Coupon payment 4. Discount 5 Market Value 6 Flotation C

B12 - X fr =B9*(1-B11) в 1000 0.09 =B2*B1 220 =B1-B4 30 1 Par Value 2 Coupon rate 3 Coupon payment 4 Discount 5 Market Value